OpenStudy (anonymous):

Completely factor the expression 112x^2-567y^2 I have multiple answers But would like to see what the right way is, so could you work it out and show your work? Thank you.

OpenStudy (anonymous):

\[112x^2-567y^2\] \[=7(16x^2-81y^2)\] \[=7(4x-9y)(4x+9y)\]